OpenStudy (anonymous):

Calculate the number of arrangements of the letters in the word NUMBER? One of the arrangement of the word NUMBER above is chosen at random.What is the probability that it begins and ends with a vowel?

OpenStudy (anonymous):

First question is easy. 6! = 720. Second question i got my answer to be: 2 x 4! = 48. Therefore probability is 48/720 = 1/15. But the answer in the book says 1/45.

OpenStudy (anonymous):

your working is right as far as i can tell...

OpenStudy (anonymous):

I'm thinking it's an error in the book?

OpenStudy (anonymous):

has to be.

OpenStudy (anonymous):

I am getting the same result as you.

OpenStudy (anonymous):

Ok thanks guys. I think it's book's mistake.

OpenStudy (amistre64):

u (4*3*2*1) e = 24 ways = 48 ways e (4*3*2*1) u = 24 ways 48 ------ = 2/30 = 1/15 ..... so, yeah 24*30